In a rectangular coordinate system a positive point charge q = +6.00 nC is placed at point x = +1.50 m, y = 0 m, and an identical charge is placed at x = -1.50 m, y = 0 m, Find the x- and y-components, the magnitude, and the direction of the electric field at x = 0, y = 0.200 m.


1
Expert's answer
2022-04-14T15:46:21-0400

"E=2E_x=2E\cos \arctg\frac yx=\frac{2kq x}{\sqrt{(x^2+y^2)^3}}" to the left.